if the vertex and focus of parabola are (3,6) and (4,5) respectively then the equation of its directrix is ????
please give solution........

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
as vertex is the mid point of foot of directrix and focus
 
hence
foot of directrix = (2, 7)
 
Now directrix is perpendicular to the line joining the focus and vertex
 
hence slope of directrix = – 1/ ( – 1) = 1
 
hence equation of directrix is
 
y – 7 = 1 ( x – 2)
 
y = x + 5